OpenStudy (anonymous):

Fan and Medal! The following function represents the profit P(n), in dollars, that a concert promoter makes by selling tickets for n dollars each: P(n) = -250n^2 + 2,500n - 4,000 Part A: What are the zeroes of the above function, and what do they represent? Show your work. (4 points) Part B: Find the maximum profit by completing the square of the function P(n). Show the steps of your work. (4 points) Part C: What is the axis of symmetry of the function P(n)? (2 points)

OpenStudy (anonymous):

\[P(n) = -250n^2 + 2,500n - 4,000 \] you want the zeros?

OpenStudy (anonymous):

I already finished part A I know the 0's are 2 and 8

OpenStudy (anonymous):

no that is wrong

OpenStudy (anonymous):

\(n\) is the price i guess if he prices the tickets at $2 or at $8 he makes no money

OpenStudy (anonymous):

The zeroes of the function would be 2 and 8. The zeroes of the function above represents the specific prices, 2 and 8, which will result in a zero profit.

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

yes looks good to me

OpenStudy (anonymous):

Okay (:

OpenStudy (anonymous):

Now how do I "complete the square"

OpenStudy (anonymous):

there is a much easier way to do it (find the max) than completing the square, but we can do it that way too if you have to

OpenStudy (anonymous):

first of all it is pretty obvious that if it is zero at 2 and at 8 then it is the biggest half way between at 5

OpenStudy (anonymous):

plus the first coordinate of the vertex is always \[-\frac{b}{2a}\] which in your case is \[-\frac{2500}{2\times (-250)}=5\]

OpenStudy (anonymous):

Yeah, but it specifically asks to complete the square (which I hate because it's so confusing)

OpenStudy (anonymous):

but we can still complete the square if you are dying to do it

OpenStudy (anonymous):

factor out the \(-250\) from the first two terms

OpenStudy (anonymous):

\[ -250n^2 + 2,500n - 4,000 \] \[-250(n^2-10n)-4,000\]

OpenStudy (anonymous):

P(n)=-250(n^2-10n)-4000?

OpenStudy (anonymous):

Okay i got that aha

OpenStudy (anonymous):

right that is step one

OpenStudy (anonymous):

then half of \(-10\) is \(-5\) complete the square via \[-250(n-5)^2-4000+25\times 250\]

OpenStudy (anonymous):

yeah i can see why

OpenStudy (anonymous):

when you replace \(n^2-10n\) by \(n-5)^2\) you have changed it

OpenStudy (anonymous):

\[(n-5)^2=n^2-10n+25\] so you have added \(25\) when you make the change

OpenStudy (anonymous):

Ohh, okay I see that now, but like if it's so many numbers and such, my brain likes to just stop reading it

OpenStudy (anonymous):

why 25 though?

OpenStudy (anonymous):

because when you square \(n-5\) you get \((n-5)^2=n^2-10n+25\)

OpenStudy (anonymous):

Ohhh

OpenStudy (anonymous):

Okay (:

OpenStudy (anonymous):

now don't forget the \(-250\) outside the parentheses

OpenStudy (anonymous):

so actually you have subtracted \(25\times 250\) so you have to add it back

OpenStudy (anonymous):

hence \[-250(n-5)^2-4000+25\times 250\]

OpenStudy (anonymous):

but why the times symbol?

OpenStudy (anonymous):

what else would you like me to write? look at this line \[-250(n-5)^2\]

OpenStudy (anonymous):

if you multiply that out you get \[-250(n^2-10n+25)\] as a first step

OpenStudy (anonymous):

I meant the times symbol for 25 x 250

OpenStudy (anonymous):

and when you multiply that out, you get \[-250n^2+2500n-6250\]

OpenStudy (anonymous):

you subtracted \(6250\) from the original expression, so you have to add it back

OpenStudy (anonymous):

there is really a much much easier way to do it

OpenStudy (anonymous):

Yeah, but my algebra teachers just love making us suffer haha

OpenStudy (anonymous):

since you know that the first coordinate of the vertex is \(5\) if you want the second coordinate, all you have to do is find \(P(5)\)

OpenStudy (anonymous):

you can pretend to to it the other way, but then just find \(P(5)\) and write that out at the end

OpenStudy (anonymous):

so after P(n)=-250(n-5)^2-4000+25*250 ,what do I do, or is that the final step?

OpenStudy (anonymous):

algebra teachers lead sad lives, most are either frustrated mathematicians or are teaching at the outer edge of their knowledge

OpenStudy (anonymous):

compute the number at the end

OpenStudy (anonymous):

-250(n-5)^2-4000+6250

OpenStudy (anonymous):

\[-400+25\times 250=2250\]

OpenStudy (anonymous):

-250(n-5)^2+2250

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And now? do I make it (n-5)(n-5)?

OpenStudy (anonymous):

@perl